OpenStudy (anonymous):

Find the point on the terminal side of θ = -3pi/4 that has an x coordinate of -1.

hero (hero):

-3pi/4 = -3(180)/4 = -3(45) = -135 = 225 Therefore, the point is (-1,-1)

OpenStudy (anonymous):

I do not understand the formula that you used, how did you get -1 as another point?

hero (hero):